OpenStudy (anonymous):

Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden? A. 16 m B. 15 m C. 14 m D. 13 m

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (snivvles.com):

16*12=192

OpenStudy (anonymous):

so its 16 then

OpenStudy (anonymous):

(9+x)*(13+x)=192 x=3 12*16=192

OpenStudy (anonymous):

Yup! c:

OpenStudy (snivvles.com):

yeah

OpenStudy (anonymous):

yeaaah